OpenStudy (anonymous):

I know the answer is 499, but cannot figure out how to get there: Find the following sum:81+64+54+48+36+36....

OpenStudy (amistre64):

does it means the mirror image?

OpenStudy (anonymous):

Or are you looking for a sequence?

OpenStudy (anonymous):

sequence. It's a question from the 2001 Math Team problem set

OpenStudy (amistre64):

86543, 1412, 26(0) 14486, 167, 28 -------- 288 might not be my idea

OpenStudy (amistre64):

to get a 9 youd need a .5 from that end and it aint got that

OpenStudy (amistre64):

the ... presupposes it goes to infinity

OpenStudy (amistre64):

81 64 54 48 36 36 17 10 6 12 0 7 4 6 12 3 2 6 1 4 3 differences dont seem to help

OpenStudy (amistre64):

81 64 54 48 36 36 9(9) 8(8) 9(6) 4(12) 6(6) 9(4) 4(16) 2(24) 9(7)+1 9(5)+3 hmmm

OpenStudy (anonymous):

I got it: \[\sum_{n=1}^{\infty}2^{n-1}3^{5-n}+\sum_{n=1}^{\infty}3^{n-1}2^{8-2n}\]

OpenStudy (anonymous):

which if you evaluate is 499

OpenStudy (anonymous):

Thanks for all the help guys! You got us thinking along the tracks we needed (my math team was looking at this problem and we couldn't figure out why the answer was 499)